A valószínűségszámítás dolgozat eredménye

1.  2.  3.  4.  5.  6.

diákok:

Faragó Anikó  ±  +  ±  +  ±  ±
Husiová Anikó  ±  ±  −  +  +  +
Juhász Árpád  ±  + ±, −  −  −  −
Kovács Szilvia  ±  +  ±  +  +  +
Lörincz Tamás  ±  ± ±, −  +  +  +
Pajkos Gabriella  +  +  ±  +  +  +
Papp Gábor János  +  +  +  +  +  +
Szabó Csilla Edit  ±  +  ±  +  +  −
Szűcs Tamás Gábor  ±  − ±, −  +  ±  +
Vadász Péter  ±  ±  −  +  +  ±
Vadnai Zoltán  −  ±  −  ±  ±  −